### v4

@@ E.2.1 Simulation design
[…]
@@
Intuitively, one may think of the considered mean-shifted contamination as careless respondents who consider only certain response patterns irrespective of the item content. Some may be straighliners after recoding of negatively keyed items: an inflated cell (5,1)(5,1)( 5 , 1 ) corresponds to recoded responses of straightliners who exclusively chose the 1st or 5th response category, (4,2)(4,2)( 4 , 2 ) to straightlining at the 2nd or 4th category, and (3,3)(3,3)( 3 , 3 ) to straightlining at the central response category.
@@ E.3.1 Simulation design
Figure E.3: Illustrative examples of the data generating process with different correlation coefficients $\rho_{*}$ (rows). The gray dots represent random draws of $(\xi,\eta)$ from the polychoric model with correlation $\rho_{*}$, whereas orange dots represent draws from the contamination distribution with correlation $-\rho_{*}$. The contamination fraction is $\varepsilon=0.15$ here, and we generate $N=10,000$ data points for a clearer visualization. The blue lines indicate the locations of the thresholds.
@@ E.3.2 Results
Finally, it is no coincidence that for all three choices of the correlation coefficient $\rho_{*}$, the three estimators arrive at a bias of about $-\rho_{*}$ at contamination level $\varepsilon=0.49$, which indicates that a near-zero correlation is estimated. Indeed, we have two groups of observations of about equal size that are symmetric to each other about the origin, one with correlation $\rho_{*}$ and one with correlation $-\rho_{*}$. It is not surprising——due to the overlap between the groups even for the robust estimator—that the best fit can be attained at a correlation estimate around 0.
@@ E.3.3 Implications for the empirical application
In our simulation with true correlation $\rho_{*}=0.9$ and contamination fraction $\varepsilon=0.15$, we obtain average correlation estimates of just below 0.9 for the robust estimator, just above 0.6 for the MLE, and just below 0.6 for the sample correlation (see Figure E.4). In the empirical application, on the other hand, we obtain estimates of 0.925-0.925- 0.925 with the robust estimator, 0.618-0.618- 0.618 with the MLE, and 0.562-0.562- 0.562 with the Pearson sample correlation (see Table 3). The robust estimator thereby assigns high Pearson residuals to the response cells $(x,y)\in\{(1,1),(1,2),(2,1),(2,2),(4,4),(4,5),(5,4),(5,5)\}$, and the empirical frequencies of these cells sum up to 11.4% (see Table 4). As discussed in Section 7.2, these cells likely correspond to careless respondents, as it is contradictory that both adjectives or neither adjective describe their personality accurately (cf. Arias et al.,, 2020). It seems plausible that these respondents simply overlooked the negation and responded to the item “envious” twice, in which case the correlation within this group is sign-flipped compared to the correlation within attentive respondents. Assuming that similar careless respondents also choose the middle response categories on both items, an overall prevalence of careless respondents around 15% seems plausible, too.
@@ Appendix F Additional results for the empirical application
[…]
1 9,780,473,685.31 15,982.37 10.81 0.14 0.35-0.35- 0.35
2 2,419.15 9.06 0.20-0.20- 0.20 0.10-0.10- 0.10 0.42
3 4.48 0.35-0.35- 0.35 0.01-0.01- 0.01 0.20-0.20- 0.20 76.11
4 0.12-0.12- 0.12 0.08-0.08- 0.08 0.39-0.39- 0.39 11.66 222,240.11
5 0.11-0.11- 0.11 0.12-0.12- 0.12 34.98 55,329.21 991,790,294,422.36
[…]
Figure F.4: Dot plot of cells for the neuroticism item adjective pair “not envious” and “envious” in the data of Arias et al., (2020), where each item has five Likert-type response options, anchored by “very inaccurate” (= 1) and “very accurate” (= 5). Each dot’s size is proportional to the relative empirical frequency of its associated contingency table cell, $\widehat{f}_{N}(x,y)$, whereas its color varies by the value of the cell’s PR, $\widehat{f}_{N}(x,y)/p_{xy}(\widehat{\bm{\theta}}_{N})-1$, at robust parameter estimate with tuning constant $c=0.6$. Note that some cells have an extremely large PR (substantially larger than the ideal value 0; see also Table F.5), indicating a poor model fit for those cells. We therefore censored the color scale so that a fully saturated red is assigned to all cells with PR 3$\geq$ 3≥ 3, i.e., the empirical frequency $\widehat{f}_{N}(x,y)$ being at least four times as large as the model frequency $p_{xy}(\widehat{\bm{\theta}}_{N})$.

### v5

@@ E.2.1 Simulation design
Figure E.1 visualizes an example dataset, which illustrates that the considered mean shifts result in negative leverage points with increasing leverage. For instance, the choice $\bm{\mu}_{H}=(0,0)^{\top}$ primarily inflates the center cell $(x,y)=(3,3)$ after discretization, $\bm{\mu}_{H}=(1,-1)^{\top}$ the cell $(x,y)=(4,2)$, and $\bm{\mu}_{H}=(2.5,-2.5)^{\top}$ the cell $(x,y)=(5,1)$. As such, these choices of mean vectors gradually move the contaminated data points (orange point clouds in Figure E.1 away from the origin. We expect that the stronger the leverage, the more bias the non-robust estimators will incur (cf. Welz et al.,, 2024). Hence, the added value of our proposed estimator should gradually increase as $\bm{\mu}_{H}$ moves away from the origin.
@@
Intuitively, one may think of the considered mean-shifted contamination as careless respondents who consider only certain response patterns irrespective of the item content. Some may be straighliners after recoding of negatively keyed items: an inflated cell $(5,1)$ corresponds to recoded responses of straightliners who exclusively chose the 1st or 5th response category, $(4,2)$ to straightlining at the 2nd or 4th category, and $(3,3)$ to straightlining at the central response category.
@@ E.3.1 Simulation design
Figure E.3: Illustrative examples of the data generating process with different correlation coefficients $\rho_{*}$ (rows). The gray dots represent random draws of $(\xi,\eta)$ from the polychoric model with correlation $\rho_{*}$, whereas orange dots represent draws from the contamination distribution with correlation $\rho_{*}$. The contamination fraction is $\varepsilon=0.15$ here, and we generate $N=10,000$ data points for a clearer visualization. The blue lines indicate the locations of the thresholds.
@@ E.3.2 Results
Finally, it is no coincidence that for all three choices of the correlation coefficient $\rho_{*}$, the three estimators arrive at a bias of about $\rho_{*}$ at contamination level $\varepsilon=0.49$, which indicates that a near-zero correlation is estimated. Indeed, we have two groups of observations of about equal size that are symmetric to each other about the origin, one with correlation $\rho_{*}$ and one with correlation $\rho_{*}$. It is not surprising——due to the overlap between the groups even for the robust estimator—that the best fit can be attained at a correlation estimate around 0.
@@ E.3.3 Implications for the empirical application
In our simulation with true correlation $\rho_{*}=0.9$ and contamination fraction $\varepsilon=0.15$, we obtain average correlation estimates of just below 0.9 for the robust estimator, just above 0.6 for the MLE, and just below 0.6 for the sample correlation (see Figure E.4). In the empirical application, on the other hand, we obtain estimates of 0.925-0.925 with the robust estimator, 0.618-0.618 with the MLE, and 0.562-0.562 with the Pearson sample correlation (see Table 3). The robust estimator thereby assigns high Pearson residuals to the response cells $(x,y)\in\{(1,1),(1,2),(2,1),(2,2),(4,4),(4,5),(5,4),(5,5)\}$, and the empirical frequencies of these cells sum up to 11.4% (see Table 4). As discussed in Section 7.2, these cells likely correspond to careless respondents, as it is contradictory that both adjectives or neither adjective describe their personality accurately (cf. Arias et al.,, 2020). It seems plausible that these respondents simply overlooked the negation and responded to the item “envious” twice, in which case the correlation within this group is sign-flipped compared to the correlation within attentive respondents. Assuming that similar careless respondents also choose the middle response categories on both items, an overall prevalence of careless respondents around 15% seems plausible, too.
@@ Appendix F Additional results for the empirical application
$X\YY$ 1 2 3 4 5
1 9,780,473,685.31 15,982.37 10.81 0.14 0.35-0.35
2 2,419.15 9.06 0.20-0.20 0.10-0.10 0.42
3 4.48 0.35-0.35 0.01-0.01 0.20-0.20 76.11
4 0.12-0.12 0.08-0.08 0.39-0.39 11.66 222,240.11
5 0.11-0.11 0.12-0.12 34.98 55,329.21 991,790,294,422.36
Table F.5: Pearson residual, $\widehat{f}_{N}(x,y)\big/p_{xy}(\widehat{\bm{\theta}}_{N})-1$, of each response $(x,y)$ for the item pair “not envious” ($X$) and “envious” ($Y$) in the measurements of Arias et al., (2020) of the neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed via the robust estimator with tuning constant $c=0.6$.
Figure F.4: Dot plot of cells for the neuroticism item adjective pair “not envious” and “envious” in the data of Arias et al., (2020), where each item has five Likert-type response options, anchored by “very inaccurate” (= 1) and “very accurate” (= 5). Each dot’s size is proportional to the relative empirical frequency of its associated contingency table cell, $\widehat{f}_{N}(x,y)$, whereas its color varies by the value of the cell’s PR, $\widehat{f}_{N}(x,y)/p_{xy}(\widehat{\bm{\theta}}_{N})-1$, at robust parameter estimate with tuning constant $c=0.6$. Note that some cells have an extremely large PR (substantially larger than the ideal value 0; see also Table F.5), indicating a poor model fit for those cells. We therefore censored the color scale so that a fully saturated red is assigned to all cells with PR $\geq 3$, i.e., the empirical frequency $\widehat{f}_{N}(x,y)$ being at least four times as large as the model frequency $p_{xy}(\widehat{\bm{\theta}}_{N})$.